# Invisible Languages of the LLM Universe

### v1

@@ Abstract
Large Language Models are trained on massive multilingual corpora, yet this abundance masks a profound crisis: of the world’s 7,613 living languages, approximately 2,000 languages with millions of speakers remain effectively invisible in digital ecosystems. We propose a critical framework connecting empirical measurements of language vitality (real-world demographic strength) and digitality (online presence) with postcolonial theory and epistemic injustice to explain why linguistic inequality in AI systems is not incidental but structural. Analyzing data across all documented human languages, we identify four categories: Strongholds (33%, high vitality and digitality), Digital Echoes (6%, high digitality despite declining vitality), Fading Voices (36%, low on both dimensions), and critically, Invisible Giants (27%, high vitality but near-zero digitality) – languages spoken by millions yet absent from LLM training data. We demonstrate that these patterns reflect continuities from colonial-era linguistic hierarchies to contemporary AI development, constituting what we term digital-epistemic injustice. Our analysis reveals that English dominance in AI is not a technical necessity but an artifact of power structures that systematically exclude marginalized linguistic knowledge. We conclude with implications for decolonizing language technology and democratizing access to AI benefits.
@@ 1 Introduction
An inquiry into content availability in Javanese, a language with 68 million speakers, yields negligible results on platforms integral to Large Language Model training. In contrast, Icelandic, spoken by 350,000 individuals, possesses a substantial digital footprint. This paradox highlights a critical issue in the manner artificial intelligence encapsulates linguistic diversity among humans. Notably, the preceding five years have observed a significant expansion of Large Language Models (LLMs), with systems such as GPT-4, Claude, and Gemini demonstrating human-like proficiency across numerous tasks (OpenAI,, 2023; Anthropic,, 2024). These models are trained utilizing extensive multilingual corpora sourced from the internet, including Common Crawl’s 159 billion web pages, Wikipedia’s 64 million articles, and vast digital archives (Gao et al.,, 2021). Nonetheless, this apparent wealth of data conceals a profound bias: a majority of the training data is derived from fewer than 20 high-resource languages, while the majority of the world’s 7,613 extant languages remain largely neglected, effectively rendering them ”invisible” (Joshi et al.,, 2020).
Prior work has documented this crisis through multiple lenses. Kornai, (2013) established that only 5% of languages will achieve meaningful digital vitality, coining the term “digital language death” to describe extinction in online spaces independent of offline endangerment. Bender et al., (2021) demonstrated that LLMs encode hegemonic worldviews, with harms falling disproportionately on marginalized communities. Simons et al., (2022) developed automated methods to measure digital language support globally, finding only 33 languages (0.4%) achieve full digital capabilities across 143 platforms.
@@
However, existing research treats linguistic exclusion primarily as a technical resource scarcity problem – languages are “low-resource” because they lack sufficient training data. This framing obscures deeper questions: Why do some languages with millions of speakers have virtually no digital presence? What historical and structural forces produce this inequality? How does the vitality-digitality gap reflect and reproduce colonial power relations?
We address these questions by integrating empirical measurement with critical theory. Building on the Invisible Lab’s comprehensive dataset measuring vitality and digitality across all 7,613 documented languages, we apply postcolonial linguistics (Said,, 1978), epistemic injustice theory (Fricker,, 2007; Helm et al.,, 2024), and critical perspectives on linguistic imperialism (Phillipson,, 1992) to explain why certain languages remain invisible despite demographic strength.
Our analysis makes three primary contributions:
First, we demonstrate that the vitality-digitality gap is not random but systematically disadvantages languages of formerly colonized regions. Invisible Giants – languages with high vitality but near-zero digitality – are concentrated in Africa, South Asia, Southeast Asia, and indigenous Americas, mirroring colonial subjugation patterns. This constitutes what we term digital-epistemic injustice: the systematic exclusion of marginalized linguistic communities from AI-mediated knowledge production.
Second, we trace continuities from colonial linguistic practices to contemporary LLM development. Missionary linguistics constructed languages as discrete, extractable objects; ISO coding systems treated them as fixed entities; contemporary NLP inherits these epistemologies. The “low-resource” framing itself reflects colonialist logics – framing African and Asian languages as deficient rather than interrogating why digital infrastructure was never built for them.
Third, we provide actionable implications for decolonizing language technology. Rather than simply “adding more data” for underrepresented languages, we argue for fundamentally reimagining LLM development: community-controlled datasets, alternative evaluation metrics centered on non-English linguistic features, and redistribution of AI economic benefits to language communities providing training data.
The urgency is clear: as LLMs become infrastructure for education, commerce, and governance, digital language vitality increasingly determines whether languages survive at all. If current trajectories continue, AI systems will accelerate rather than reverse language endangerment, creating feedback loops where digital invisibility causes offline decline.
@@ 2.1 Digital Language Death and the Vitality-Digitality Distinction
Kornai, (2013) established a crucial insight: digital vitality operates independently from traditional language vitality. A language can remain spoken intergenerationally (high vitality by EGIDS standards) yet be functionally extinct online (zero digitality). Using machine learning on 8,426 languages, Kornai created a four-tier classification finding that languages relegated to “Heritage” status – digitally archived but not used for living communication – cannot digitally ascend regardless of documentation efforts.
This distinction reveals a fundamental gap in endangerment frameworks. The Expanded Graded Intergenerational Disruption Scale (EGIDS) (Lewis et al.,, 2010) assesses offline vitality through intergenerational transmission and institutional use, classifying languages from 0 (International) to 10 (Extinct). UNESCO’s nine-factor framework similarly focuses on speaker attitudes, policies, and documentation (on Endangered Languages,, 2003). Yet these frameworks emerged before the internet became central to economic and social participation.
Recent work has attempted to bridge this gap. The Digital Language Diversity Project developed a six-level Digital Language Vitality Scale assessing capacity, presence, and performance across digital domains (Soria and Ballatore,, 2017). Simons et al., (2022) introduced automated measurement across seven support categories (content, encoding, translation, analysis) and 143 platforms. However, no framework successfully integrates traditional and digital vitality into a unified model validated across diverse linguistic contexts.
We adopt a two-dimensional framework treating vitality and digitality as independent axes. Vitality measures socio-demographic presence: first-language speaker volume (from Ethnologue) and EGIDS status, combined through unsupervised clustering. Digitality measures online footprint: prevalence across Common Crawl (159 billion pages), Wikipedia (64 million articles), Hugging Face datasets (114,000) and models (447,000), and open language archives (474,000 entries). The representation score equals digitality minus vitality, quantifying the gap between real-world strength and digital presence.
@@ 2.2 Epistemic Injustice in Language Technology
Helm et al., (2024) introduced “language modeling bias” as by-design preferences for certain languages embedded in AI systems, arguing this constitutes epistemic injustice – systems are precise for dominant powers but limited in expressing socio-culturally relevant concepts of marginalized communities. This framework, drawing on Fricker, (2007), distinguishes testimonial injustice (dismissing contributions based on identity prejudice) from hermeneutical injustice (gaps in collective knowledge resources enabling interpretation of social experiences).
LLMs contribute to both forms at unprecedented scale. Testimonial injustice occurs when models systematically downrank, mistranslate, or fail to recognize contributions in non-dominant languages. When a Swahili speaker’s input receives degraded performance compared to English, the system effectively judges their testimony as less credible. Hermeneutical injustice occurs when languages lack vocabulary in model training data to express community-specific concepts – there are no isiZulu words for “dinosaur” or “evolution” in most LLM training sets, making scientific discourse in isiZulu impossible (Nekoto et al.,, 2020).
Recent work extends this framework. Talat et al., (2024) argue that labeling certain data as “knowledge” in knowledge-enhanced LLMs obscures harm to marginalized groups, as knowledge enhancement perpetuates epistemic injustice without diversification. Immigration systems relying on limited machine translation endanger lives when asylum seekers cannot articulate claims in languages systems recognize (Garibay and González,, 2021). The opacity and scale of AI systems facilitate epistemic injustice at levels previously impossible – billions of interactions daily embed the message that certain languages and ways of knowing don’t matter.
We extend epistemic injustice theory by connecting it to digital language vitality. The vitality-digitality gap directly produces hermeneutical injustice: communities with rich oral traditions and millions of speakers cannot access AI benefits because their linguistic resources were never digitized. This is not accidental scarcity but structural exclusion reflecting historical marginalization.
@@ 2.3 Postcolonial Linguistics and Digital Imperialism
Said, (1978) defined orientalism as ways Western cultures dominate and reconstruct Eastern cultures, depicting them as undeveloped and inferior. This framework illuminates how linguistic hierarchies established during colonialism persist in AI development. English dominance is not a neutral technical outcome but reflects British imperial expansion and subsequent American technological hegemony.
Contemporary language technology exhibits what Mohamed et al., (2020) term “AI colonialism” – extractive relationships where Global North institutions mine data from Global South communities, build models on that data, then sell tools back to those communities at profit. This mirrors classic colonial resource extraction. South African facial recognition creating digital apartheid (Benjamin,, 2023), Venezuelan data labeling under exploitative conditions (Gray and Suri,, 2019), and Māori communities fighting for data sovereignty (Kukutai and Taylor,, 2016) exemplify continuities from territorial colonialism to data colonialism.
The “low-resource language” framing itself embodies colonialist logics. As Dotan et al., (2023) argue, African languages remain “low resource” not due to inherent deficiency but because Western institutions never invested in digital infrastructure for them. The term positions these languages as lacking, requiring benevolent intervention from well-resourced institutions, rather than recognizing that resource scarcity is a political product of marginalization.
Daems, (2024) trace material continuities: missionary linguistics’ construction of languages as discrete, extractable objects prefigures modern NLP’s treatment of languages as datasets to be harvested. ISO 639 language codes, developed by Western institutions, treat languages as fixed entities rather than fluid practices, enabling computational processing but erasing sociolinguistic realities. Contemporary commercial technology providers still draw on colonial-era translations and grammatical descriptions produced by missionaries.
Self-orientalization compounds these dynamics. Liu and Li, (2017) document how some Chinese professionals in Australia actively cater to stereotypes, labeling themselves with Confucian attributes to gain workplace advantage. Chen, (2021) shows tourism discourse in China Daily depicts some regions as marginalized due to lack of English service, with “Chinglish” reinforcing stereotypes of Chinese English inadequacy – a form of internalized linguistic imperialism that LLMs trained on such texts will reproduce.
@@ 2.4 Language Ideology and Technological Design
Language ideologies – beliefs about language that rationalize social structures – are embedded in technological architecture (Woolard and Schieffelin,, 1998). Hohn et al., (2024) demonstrate that conversational AI embeds Western linguistic ideologies privileging certain ways of speaking: standardized grammar over vernacular, written over oral, explicitness over context-dependence, monolingualism over multilingualism. These ideological commitments shape design choices about what counts as “good” language data, which features to optimize, and how to evaluate performance.
The ideology of “Standard Language” positions prestigious varieties as correct while stigmatizing others as deficient (Lippi-Green,, 2012). LLMs overwhelmingly train on formal written language, embedding prescriptive norms. When models correct African American Vernacular English to Standard American English, they encode racist language ideologies positioning AAVE as “broken” English (Blodgett et al.,, 2020). Similar dynamics operate multilingually – “correct” Hindi means formal written Hindi, erasing Hinglish and regional varieties actually used by millions.
Platform architectures encode linguistic assumptions. Twitter’s character limits advantage logographic writing systems over alphabetic ones. TikTok’s audio-visual primacy advantages languages with existing popular media. LinkedIn’s professional register expectations advantage languages with established business communication norms. Yet these platform-specific language ideologies remain understudied – we lack systematic analysis of how affordances advantage or disadvantage different languages and practices.

### v2

@@ Abstract
Large Language Models are trained on massive multilingual corpora, yet this abundance masks a profound crisis: of the world’s 7,613 living languages, approximately 2,000 languages with millions of speakers remain effectively invisible in digital ecosystems. We propose a critical framework connecting empirical measurements of language vitality (real-world demographic strength) and digitality (online presence) with postcolonial theory and epistemic injustice to explain why linguistic inequality in AI systems is not incidental but structural. Analyzing data across all documented human languages, we identify four categories: Strongholds (33%, high vitality and digitality), Digital Echoes (6%, high digitality despite declining vitality), Fading Voices (36%, low on both dimensions), and critically, Invisible Giants (27%, high vitality but near-zero digitality) – languages spoken by millions yet absent from LLM training data. We demonstrate that these patterns reflect continuities from colonial-era linguistic hierarchies to contemporary AI development. Our analysis reveals that English dominance in AI is not a technical necessity but an artifact of power structures that systematically exclude marginalized linguistic knowledge. We conclude with implications for decolonizing language technology and democratizing access to AI benefits.
@@ 1 Introduction
An examination of the availability of digital content in Javanese – a language with more than 69 million speakers – reveals minimal representation on platforms that are central to the training of Large Language Models (LLMs). By contrast, Icelandic, with approximately 320,000 speakers, exhibits a disproportionately large digital presence. This apparent paradox underscores a fundamental problem in the way contemporary artificial intelligence systems capture and model human linguistic diversity. Over the past five years, LLMs such as GPT-4, Claude, and Gemini have expanded rapidly, achieving performance that approximates human-level proficiency across a broad array of tasks (achiam2023gpt; anthropic2024claude). These models are trained on massive multilingual corpora harvested from the internet, including Common Crawl’s roughly 300 billion web pages, Wikipedia’s more than 64 million articles, and large-scale digital text repositories (gao2020pile). However, this ostensible abundance of data obscures a substantial structural bias: the overwhelming majority of training material originates from fewer than twenty high-resource languages, while most of the world’s 7,613 documented languages receive negligible coverage and are effectively rendered invisible in these models (eberhard2015ethnologue; joshi2020state).
Prior scholarship has characterized this crisis from multiple perspectives. kornai2013digital estimated that only approximately 5% of the world’s languages are likely to attain substantive digital vitality, and introduced the concept of “digital language death” to denote language extinction in online environments as a problem distinct from, and additional to, conventional offline endangerment. bender2021dangers showed that large language models encode and propagate hegemonic worldviews, with the attendant harms disproportionately affecting marginalized communities. Complementarily, simons2022assessing proposed automated methodologies to quantify digital language support at a global scale, reporting that only 33 languages (0.4%) exhibit full digital functionality across 143 surveyed platforms.
@@
However, the predominant body of research conceptualizes linguistic exclusion primarily as a problem of technical resource scarcity, wherein languages are classified as ‘low-resource’ on the basis of insufficient training data. This framing, however, obscures more fundamental lines of inquiry: why do some languages with millions of speakers possess virtually no digital footprint? Which historical, political, and structural dynamics give rise to this inequality? Moreover, how does the divergence between on-the-ground prevalence (or vitality) and online or computational prevalence (or digitality) both reflect and reproduce colonial and neocolonial power relations?
We address these research questions through an integration of large-scale empirical measurement and critical theoretical frameworks. Drawing on a comprehensive dataset that quantifies linguistic vitality and digital presence for all 7,613 documented languages, we build postcolonial linguistic theories from said1977orientalism and levisen2019postcolonial, conceptualizations of epistemic injustice (fricker2007epistemic; helm2024diversity), and critical analyses of linguistic imperialism (phillipson1992linguistic). This combined approach enables us to investigate the underlying structural, ideological, and historical mechanisms that explain why certain languages remain marginalized or rendered invisible in digital and epistemic spaces despite substantial demographic strength.
Our analysis yields three primary contributions. First, we demonstrate that the vitality-digitality gap is not neutral; rather, it systematically disadvantages languages spoken in formerly colonized regions. Invisible Giants – languages exhibiting high sociolinguistic vitality but near-zero digital presence – are disproportionately concentrated in Africa, South Asia, Southeast Asia, and Indigenous regions of the Americas, thereby recapitulating historical patterns of colonial domination. We conceptualize this as a form of digital epistemic injustice, namely, the systematic exclusion of marginalized linguistic communities from AI-mediated processes of knowledge production. Second, we trace continuities from colonial-era linguistic practices to contemporary large language model (LLM) development. Missionary linguistics frequently construed languages as discrete, extractable objects; subsequent ISO coding regimes further reified them as fixed and enumerable entities; contemporary natural language processing (NLP) methodologies inherit and operationalize these epistemic assumptions. The prevailing ‘low-resource’ characterization itself reproduces colonial logics, positioning African and Asian languages as inherently deficient rather than interrogating the political-economic conditions under which digital and computational infrastructures were never developed for them. Third, we articulate concrete implications for the decolonization of language technologies. In contrast to approaches that focus primarily on ‘adding more data’ for underrepresented languages, we argue for a fundamental reconfiguration of LLM development practices. This includes the use of community-governed datasets, the design of alternative evaluation metrics that prioritize contextualized, non-majority linguistic features, and mechanisms for redistributing the economic value generated by AI systems to the language communities whose data underpins model training.
The situation presents a clear urgency: as large language models increasingly function as core infrastructure for education, commerce, and governance, the vitality of languages in digital environments is becoming a critical determinant of their long-term survival. If current development trajectories persist, artificial intelligence systems are more likely to intensify than mitigate processes of language endangerment, thereby establishing self-reinforcing feedback loops in which digital marginalization contributes directly to accelerated offline linguistic decline.
@@ 2.1 Digital Language Death and the Vitality-Digitality Distinction
lewis2010assessing introduced the EGIDS (Expanded Graded Intergenerational Disruption Scale) as a systematic framework for assessing language vitality and endangerment, extending the original GIDS (Graded Intergenerational Disruption Scale) proposed by fishman1991reversing. Building on this line of inquiry, kornai2013digital demonstrated that digital vitality constitutes an autonomous dimension, largely decoupled from traditional measures of language vitality. Thus, a language may exhibit robust intergenerational transmission and therefore rank highly on EGIDS, while simultaneously being effectively absent from the digital sphere (i.e., exhibiting negligible or zero digital presence). Drawing on a large-scale dataset comprising 8,426 languages, Kornai proposed a four-tier typology of digital status and showed that languages classified as heritage – those that are digitally documented but no longer employed in active, everyday communication – have a very low probability of achieving higher levels of digital presence, irrespective of the extent of documentation efforts.
This distinction exposes a fundamental limitation in existing language endangerment assessment frameworks. The EGIDS evaluates offline language vitality primarily on the basis of intergenerational transmission and domains of institutional use, assigning languages to categories ranging from 0 (International) to 10 (Extinct). Similarly, UNESCO’s six-level framework emphasizes dimensions such as speaker attitudes, language policies, and the extent of linguistic documentation (unesco2003language). However, both frameworks were envisioned prior to the point at which the internet became a central medium for economic activity and social participation.
Recent scholarship has sought to address this discrepancy. The Digital Language Diversity Project proposed a six-level language vitality scale that evaluates capacity, presence, and performance across multiple digital domains (soria2018digital). In parallel, simons2022assessing introduced an automated evaluation framework encompassing seven support categories (from content presence to modern speech assistance) and 143 technological platforms. Nonetheless, no existing framework has yet succeeded in systematically integrating grassroots and digital dimensions of language presence and strength into a coherent model that has been empirically validated across a broad range of linguistic contexts.
We employ a two-dimensional analytical framework in which vitality and digitality are modeled as orthogonal axes. Vitality captures the strength of a language in grassroots terms, operationalized via the estimated number of first-language speakers (sourced from Ethnologue) and its current EGIDS classification. Digitality quantifies the language’s presence in digital environments, measured through its representation in Common Crawl (159 billion web pages analyzed), Wikipedia (64 million articles), Hugging Face datasets (114,000) and models (447,000), as well as open language archives (474,000 entries). For both vitality and digitality, we aggregate the respective indicators using factor analysis to obtain single standardized vitality and digitality scores. On this basis, we define a representation score as the difference between the digitality and vitality scores, which serves as an approximation of the discrepancy between a language’s sociolinguistic strength and its digital representation.
@@ 2.2 Epistemic Injustice in Language Technology
The notion of language modeling bias introduced by helm2024diversity denotes built-in preferences for particular languages within AI systems. They argue that this bias constitutes a form of epistemic injustice: such systems tend to exhibit high precision for languages associated with dominant social groups, while being substantially less capable of representing socio-culturally salient concepts specific to marginalized communities. Large language models (LLMs) thus have the potential to reproduce and amplify both testimonial injustice (the unjust deflation of a speaker’s credibility due to identity prejudice) and hermeneutical injustice (structural gaps in shared interpretive resources that hinder the understanding of certain social experiences) at an unprecedented scale. Testimonial injustice manifests when models systematically downrank, mistranslate, or fail to recognize contributions articulated in non-dominant languages. For example, an input in Swahili receives systematically lower-quality processing than an equivalent input in English, such that the system effectively treats the Swahili speaker’s testimony as less reliable or less worthy of engagement. Hermenutical injustice occurs when the training data for a given language lack the lexical or conceptual resources necessary to articulate community-specific or domain-specific concepts. For instance, if isiZulu training data do not contain terms corresponding to dinosaurs or evolution, the LLM’s ability to support, extend, or even enable scientific discourse in isiZulu is significantly constrained (nekoto2020participatory).
Recent scholarship extends this framework by arguing that designating particular data as ‘knowledge’ within knowledge-enhanced large language models (LLMs) can obscure harms experienced by marginalized groups. In the absence of meaningful diversification, such knowledge-enhancement practices risk reproducing and amplifying epistemic injustice. For example, immigration and asylum adjudication processes that rely on narrowly scoped machine translation systems may place lives at risk when asylum seekers are unable to articulate their claims in languages that these systems recognize or adequately process. The opacity, scale, and infrastructural embeddedness of contemporary AI systems enable epistemic injustice on an unprecedented magnitude: billions of interactions per day can implicitly communicate that certain languages, epistemic traditions, and modes of knowing are inconsequential or invalid.
We extend this line of inquiry by situating it within the framework of digital language vitality. The discrepancy between a language’s sociolinguistic vitality and its level of digital representation constitutes a direct source of hermeneutical injustice. Specifically, communities with robust oral traditions and large speaker populations are systematically excluded from the benefits of artificial intelligence because their linguistic resources have not been digitized or incorporated into computational infrastructures. This condition does not reflect incidental scarcity, but rather a form of structural exclusion rooted in, and perpetuating, historical patterns of marginalization.
@@ 2.3 Postcolonial Linguistics and Digital Imperialism
said1977orientalism conceptualized Orientalism as a discursive formation through which Western societies exert dominance over, and systematically reconfigure, Eastern societies, representing them as underdeveloped and inferior. Building on this, ismail2024revisiting demonstrate how linguistic hierarchies established during colonial rule continue to shape the development of modern technologies. The predominance of English in AI systems is thus not a neutral or purely technical outcome, but rather a manifestation of historical imperial expansion and the ensuing technological hegemonies that structure contemporary knowledge production.
Contemporary language technologies exhibit extractive dynamics in which institutions located predominantly in the Global North appropriate data from communities in the Global South, develop computational models based on these data, and subsequently commercialize and resell the resulting tools to those same communities for profit (mohamed2020decolonial). Phenomena such as the production of algorithmic bias in facial recognition systems (ruha2019race), data annotation performed under exploitative labor conditions (gray2019ghost), and community-led struggles for data sovereignty (kukutai2016indigenous) illustrate multiple lines of continuity between historical forms of territorial colonialism and contemporary data colonialism. These processes reproduce classical patterns of colonial resource extraction and reinforce longstanding power asymmetries between the West and the Global South in the domain of language, now intensified and reconfigured through the deployment of artificial intelligence in the digital age.
The very designation of a ‘low-resource language’ can be understood as reproducing a colonial epistemic framework. As adebara2022towards argue, many African languages are currently classified as low resource not due to any inherent deficiency, but because Western institutions have historically failed to invest in the digital infrastructure necessary for their computational development. This terminology thereby constructs these languages as intrinsically lacking and in need of benevolent support from already well-resourced institutions, rather than foregrounding the fact that resource scarcity is a politically and historically produced outcome of systemic marginalization.
migge2025material examine historical-material continuities between early missionary linguistics and contemporary natural language processing. Missionary-era construction of ‘languages’ as discrete, extractable, and ostensibly bounded objects predates modern NLP’s conceptualization of languages as datasets to be collected, segmented, and exploited. ISO-639-3 language codes, developed within Western standardization regimes, similarly presuppose languages as stable, clearly delimited entities rather than as dynamic, heterogeneous, and inherently fluid sociolinguistic practices. This ontological framing facilitates large-scale computational processing while simultaneously obscuring and erasing the complexity of lived linguistic repertoires. Moreover, contemporary commercial language technologies continue to rely heavily on colonial-era translations and grammatical descriptions produced by missionaries, thereby reproducing and extending the epistemic and material effects of those earlier regimes of linguistic knowledge production.
Self-orientalization further intensifies these dynamics. yan2018other document how certain Chinese professionals strategically mobilize and even amplify cultural stereotypes, self identifying with Confucian traits, in order to secure advantages in professional settings. Similarly, chen2021language demonstrate that tourism discourse in China not only constructs particular regions as peripheral on the basis of limited English-language services, but also reiterates tropes of Chinese English deficiency. This discursive pattern constitutes a form of internalized linguistic imperialism that is likely to be replicated and amplified by large language models (LLMs) trained on such corpora.
@@ 2.4 Language Ideology and Technological Design
Language ideologies – belief systems concerning language that serve to legitimize and naturalize existing social structures – are necessarily instantiated within technological infrastructures (woolard2020language). Contemporary conversational AI systems operationalize Western linguistic ideologies that privilege specific modes of language use, including standardized grammar over vernacular varieties, written over oral forms, explicitness over contextual inference, and monolingualism over multilingual repertoires. These ideological commitments inform core design decisions regarding what is classified as high-quality language data, which linguistic features are prioritized for optimization, and which criteria are employed to assess system performance.
The ideology of a ‘standard language’ constructs socially prestigious varieties as normatively correct while marking other varieties as deficient or erroneous (lippi2012english). Large language models (LLMs) are trained predominantly on formal written registers, thereby internalizing and reproducing prescriptive linguistic norms. When such models ‘correct’ African American Vernacular English into Standard American English, they operationalize and reinforce raciolinguistic ideologies that frame the former as improper or broken (blodgett2020language). Analogous processes occur in multilingual contexts: what is labeled as ‘correct’ Hindi is typically formal written Hindi, which marginalizes and effectively erases Hinglish and diverse regional varieties that constitute the everyday linguistic practices of millions of speakers.
Platform architectures embed and reproduce particular linguistic assumptions. For example, Twitter’s character constraints structurally privilege logographic writing systems relative to alphabetic ones. TikTok’s emphasis on audio‑visual content tends to favor languages that already possess extensive, globally circulating popular media repertoires. LinkedIn’s normative expectations of a professional register advantage languages with long-established conventions of formal business communication. Despite their widespread influence, these platform-specific language ideologies have received limited scholarly attention; systematic analyses are still lacking regarding how platform affordances differentially advantage or disadvantage specific languages and communicative practices.